# Singular Bayesian Neural Networks

### v1

@@ E.7.1 Explicit Complexity Bound for Pinto’s Class
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\},$
@@
  • $r_{j}$ is the rank bound for layer $j$.
@@
    First layer: The term $\|W_{1}\|_{F}\sqrt{h_{1}}$ for the first layer appears because the first layer is not constrained by rank in the bound, only by spectral norm and Frobenius norm.
@@
Note on depth dependence: Both the rank-constrained bound from Theorem E.17 and standard bounds contain a product over depth, specifically the factor $\prod_{i=j}^{D}C_{0}C_{i}$ which grows as $C_{0}^{D-j+1}\prod_{i=j}^{D}C_{i}$. As noted by Pinto et al. (2025), this unfavorable exponential factor in depth may be an artifact of the current proof technique rather than a limitation. When comparing the rank-constrained bound to existing bounds, the key improvement comes from the rank-dependent terms $\sqrt{r_{j}h_{j}}$ rather than from the depth-dependent product, which appears in both settings.
@@
Substituting the explicit bound from Theorem E.17:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}.$
@@
$\quad+\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}$
@@ E.9 Comparison with PAC-Bayes Bounds
Consider a single-layer network with input dimension $d_{\text{in}}$, output dimension $d_{\text{out}}$, and rank $r$.
@@
$\text{Gap}_{\text{PAC}}=O\left(\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})}{m}}\right),$ (15)
with rank dependence $\sqrt{r}$ and uniform $\frac{1}{\sqrt{m}}$ scaling in sample size (up to logarithmic factors).
@@
$\text{Gap}_{\text{GC}}=O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)+O\left(\frac{1}{\sqrt{m}}\right),$ (16)
where:
  • The first term has rank dependence $\sqrt{r}$ and scales as $\frac{1}{m}$ (for fixed $\|X\|_{F}$);
  • The second term (concentration) scales as $\frac{1}{\sqrt{m}}$ and is rank-independent;
  • For large $m$, the $O(1/\sqrt{m})$ term dominates;
  • For moderate $m$ with large $\|X\|_{F}$, the $O(1/m)$ term may dominate.
If, in addition, inputs are uniformly bounded as $\|x_{j}\|_{2}\leq R$, then $\|X\|_{F}=O(\sqrt{m})$ and the rank-dependent term becomes $O(\sqrt{rd_{\text{out}}/m})$.
@@
$\sqrt{\frac{\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})+\log(2\sqrt{m}/\delta)}{2m}}\leq\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}}.$
Ignoring logarithmic terms and constants, this scales as
$O\left(\sqrt{\frac{r(d_{\text{in}}+d_{\text{out}})}{m}}\right)=\sqrt{r}\cdot O\left(\sqrt{\frac{d_{\text{in}}+d_{\text{out}}}{m}}\right),$
so the rank dependence is $\sqrt{r}$ and the sample-size dependence is $1/\sqrt{m}$.
@@
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})\leq\frac{\|X\|_{F}}{m}\|W_{1}\|_{F}\sqrt{h_{1}}C_{0}C_{1},$
where $h_{1}=d_{\text{out}}$. For a rank-$r$ matrix $W_{1}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ with $\|W_{1}\|_{2}\leq C_{1}$, we have
$\|W_{1}\|_{F}\leq\sqrt{r}\,\|W_{1}\|_{2}\leq\sqrt{r}C_{1}.$
Substituting with $h_{1}=d_{\text{out}}$,
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})$ $\leq\frac{\|X\|_{F}}{m}\cdot\sqrt{r}C_{1}\cdot\sqrt{d_{\text{out}}}\cdot C_{0}C_{1}$
$=\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{rd_{\text{out}}}}{m}$
$=O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right).$
@@
$\mathbb{E}[\ell(f(x),y)]$ $\leq\hat{L}(f)+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2m}}$
$=\hat{L}(f)+O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)+O\left(\frac{1}{\sqrt{m}}\right),$
where the first term is rank-dependent and decays as $1/m$ (for fixed $\|X\|_{F}$), while the second term is rank-independent and decays as $1/\sqrt{m}$. This yields the claimed form for $\text{Gap}_{\text{GC}}$.
If we additionally assume $\|x_{j}\|_{2}\leq R$ for all $j$, then
$\|X\|_{F}^{2}=\sum_{j=1}^{m}\|x_{j}\|_{2}^{2}\leq mR^{2}\quad\Rightarrow\quad\|X\|_{F}\leq R\sqrt{m},$
so the rank-dependent term simplifies to
$O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)=O\left(\frac{\sqrt{m}\sqrt{rd_{\text{out}}}}{m}\right)=O\left(\sqrt{\frac{rd_{\text{out}}}{m}}\right).$
@@
For large $m$, the slower-decaying $O(1/\sqrt{m})$ term dominates; for moderate $m$ and large $\|X\|_{F}\sqrt{rd_{\text{out}}}$, the $O(1/m)$ term can dominate.
@@ E.9.1 Complementary Strengths
    Large sample regime: For large $m$, both bounds scale as $O(1/\sqrt{m})$, but PAC-Bayes may have better constants when the posterior is concentrated (small $C_{\max}$).
@@
    Posterior concentration: When the posterior is highly concentrated around a good solution, $C_{\max}$ can be very small, making the KL term tight.
@@
    Large input norms: PAC-Bayes bounds do not depend on $\|X\|_{F}$, which is advantageous when inputs have very large norm or heavy tails.
@@
    Dimension regime: When $r(d_{\text{in}}+d_{\text{out}})$ is small relative to $\|X\|_{F}^{2}d_{\text{out}}$, the PAC-Bayes bound can be more favorable.
Remark E.24 (When Gaussian complexity is stronger).
Gaussian complexity bounds might be tighter when:
  1. 1.
    Small to moderate sample sizes: When $m$ is not too large and $\|X\|_{F}\sqrt{rd_{\text{out}}}=O(\sqrt{m})$, the $O(1/m)$ term provides faster decay than the $O(1/\sqrt{m})$ concentration term.
  2. 2.
    Normalized or bounded inputs: When $\|X\|_{F}=O(\sqrt{m})$ (e.g., inputs normalized to unit norm), the rank-dependent term becomes $O(\sqrt{rd_{\text{out}}/m})$, which can be tighter than PAC-Bayes when $rd_{\text{out}}\ll r(d_{\text{in}}+d_{\text{out}})$.
  3. 3.
    No prior specification required: Unlike PAC-Bayes, which requires choosing a prior $P$, Gaussian complexity bounds depend only on the realized network architecture and data.
  4. 4.
    Architectural transparency: The bound explicitly shows how rank, spectral norms, and layer dimensions interact through the network depth, providing architectural guidance.
  5. 5.
    Connection to deterministic theory: Gaussian complexity bounds inherit the full machinery of deterministic generalization theory, including recent advances in rank-sensitive analysis.
@@ E.10.1 Setup and Notation
Define the ”good event” $\mathcal{E}_{\text{good}}$ as the intersection of all these events across all $D$ layers:
@@ E.10.2 Approach A: Decomposition of the BNN Predictor
We can decompose this expectation by conditioning on whether $\mathcal{E}_{\text{good}}$ occurs:
$f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}]\cdot\mathbb{P}(\mathcal{E}_{\text{good}})+\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{bad}}]\cdot\mathbb{P}(\mathcal{E}_{\text{bad}}),$ (22)
where $\mathcal{E}_{\text{bad}}=\mathcal{E}_{\text{good}}^{c}$.
Analysis of the ”good” term.
Define the conditional predictor:
$f_{\mathrm{BNN}}^{\text{good}}(x;\theta):=\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}].$ (23)
This conditional expectation is taken only over samples that satisfy the spectral norm bounds. Replacing the full posterior with the conditional distribution given $\mathcal{E}_{\text{good}}$, we have by the same argument as Lemma E.11:
@@
where $\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)$ consists of network functions realized by samples in $\mathcal{E}_{\text{good}}$.
Since all such samples satisfy the spectral norm constraints by construction, we have:
@@
Therefore, by the same arguments as Proposition E.13:
$\hat{\mathcal{G}}_{S_{x}}(\{f_{\mathrm{BNN}}^{\text{good}}\})\leq\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r)).$ (26)
Analysis of the ”bad” term.
The contribution from $\mathcal{E}_{\text{bad}}$ has small probability $\delta_{\text{total}}$. If the loss function $\ell$ is bounded in $[0,1]$, this contributes at most $\delta_{\text{total}}$ to the expected loss.
@@
Let $\ell:\mathbb{R}^{k}\times\mathbb{R}^{k}\to[0,1]$ be an $\mathcal{L}$-Lipschitz loss function. The expected loss of the BNN predictor satisfies:
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $=\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)\mid\mathcal{E}_{\text{good}}]\cdot\mathbb{P}(\mathcal{E}_{\text{good}})$
$\quad+\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)\mid\mathcal{E}_{\text{bad}}]\cdot\mathbb{P}(\mathcal{E}_{\text{bad}})$
$\leq\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+1\cdot\delta_{\text{total}}.$ (27)
Applying the generalization bound from Theorem E.15 to $f_{\mathrm{BNN}}^{\text{good}}$ (following the proof of Theorem 3.9):
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{i=1}^{m}\ell(f_{\mathrm{BNN}}^{\text{good}}(x_{i}),y_{i})$
$\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}+\delta_{\text{total}}.$ (28)
@@
The additional term $\delta_{\text{total}}$ represents the ”cost” of using high-probability bounds instead of almost-sure bounds. For Gaussian posteriors with Van Handel-type bounds (see Remark E.30), choosing $\delta_{i}^{A}=\delta_{i}^{B}=\delta_{0}/(2D)$ for each layer gives $\delta_{\text{total}}=\delta_{0}$. The spectral norm bounds $C_{i}^{A}$ and $C_{i}^{B}$ then scale as $O(\sqrt{\log(D/\delta_{0})})$, which introduces a mild logarithmic dependence on the depth and confidence level.

### v2

@@ E.7.1 Explicit Complexity Bound for Pinto’s Class
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{B_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\},$
@@
  • $r_{j}$ is the rank bound for layer $j$,
  • $B_{1}:=\sup_{f_{W}\in\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)}\|W_{1}\|_{F}$ is the worst-case Frobenius norm of the first-layer weight matrix over the class. Since $\mathrm{rank}(W_{1})\leq r_{1}$ and $\|W_{1}\|_{2}\leq C_{1}$ for all members of the class, we have $B_{1}\leq\sqrt{r_{1}}\,C_{1}$.
@@
    First layer: The term $B_{1}\sqrt{h_{1}}$ for the first layer differs in form from the $\sqrt{r_{j}h_{j}}$ terms for deeper layers. This is an artifact of the layer-peeling proof technique (specifically, how the chaining argument handles the input layer), not a reflection of a missing rank constraint. The rank of the first layer still enters through $B_{1}\leq\sqrt{r_{1}}\,C_{1}$, so that the first-layer contribution is at most $\sqrt{r_{1}}\,C_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}$, which has the same $\sqrt{r_{1}}$ dependence as deeper layers.
@@
Note on depth dependence: Both the rank-constrained bound from Theorem E.17 and standard bounds contain a product over depth, specifically the factor $\prod_{i=j}^{D}C_{0}C_{i}$ which grows as $C_{0}^{D-j+1}\prod_{i=j}^{D}C_{i}$. As noted by Pinto et al. (2025), this unfavorable exponential factor in depth may be an artifact of the current proof technique rather than a fundamental limitation. When comparing the rank-constrained bound to existing bounds, the key improvement comes from the rank-dependent terms $\sqrt{r_{j}h_{j}}$ rather than from the depth-dependent product, which appears in both settings.
@@
Substituting the explicit bound from Theorem E.17, and using $B_{1}\leq\sqrt{r_{1}}\,C_{1}$:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\sqrt{r_{1}}\,C_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}.$
@@
$\quad+\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}}{m}\left\{\sqrt{r_{1}}\,C_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}$
@@ E.9 Comparison with PAC-Bayes Bounds
Consider a single-layer network with input dimension $d_{\text{in}}$, output dimension $d_{\text{out}}$, and rank $r$. Let $\ell(\cdot,y)\in[0,1]$ be $\mathcal{L}$-Lipschitz in its first argument.
@@
$\text{Gap}_{\text{PAC}}\;\leq\;\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}},$ (15)
where $C_{\max}$ is the maximum per-parameter KL divergence between the posterior and the prior. The structural scaling (suppressing logarithmic terms, $C_{\max}$, and constants) is
$\text{Gap}_{\text{PAC}}=O\!\left(\sqrt{\frac{r(d_{\text{in}}+d_{\text{out}})}{m}}\right).$
@@
$\text{Gap}_{\text{GC}}\;\leq\;\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{r\,d_{\text{out}}}}{m}\;+\;3\sqrt{\frac{\log(2/\delta)}{2m}},$ (16)
where $C_{0}$ is the universal chaining constant and $C_{1}$ is the spectral norm bound.
If inputs are uniformly bounded as $\|x_{j}\|_{2}\leq R$, then $\|X\|_{F}\leq R\sqrt{m}$ and the rank-dependent term simplifies to $O(\sqrt{r\,d_{\text{out}}/m})$.
@@
$\sqrt{\frac{\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})+\log(2\sqrt{m}/\delta)}{2m}}\leq\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}},$
which yields the expression in (15). Dropping logarithmic terms and constants gives the structural scaling $O(\sqrt{r(d_{\text{in}}+d_{\text{out}})/m})$.
@@
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})\leq\frac{\|X\|_{F}}{m}\cdot B_{1}\sqrt{h_{1}}\cdot C_{0}C_{1},$
where $h_{1}=d_{\text{out}}$. Since $B_{1}\leq\sqrt{r}\,C_{1}$ (see Theorem E.17), substituting yields
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})$ $\leq\frac{\|X\|_{F}}{m}\cdot\sqrt{r}\,C_{1}\cdot\sqrt{d_{\text{out}}}\cdot C_{0}C_{1}=\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{r\,d_{\text{out}}}}{m}.$
@@
$\text{Gap}_{\text{GC}}\leq\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{r\,d_{\text{out}}}}{m}+3\sqrt{\frac{\log(2/\delta)}{2m}},$
which is (16). The first (rank-dependent) term decays as $1/m$ for fixed $\|X\|_{F}$, while the second (concentration) term decays as $1/\sqrt{m}$.
If we additionally assume $\|x_{j}\|_{2}\leq R$ for all $j$, then $\|X\|_{F}\leq R\sqrt{m}$, so the rank-dependent term becomes
$\sqrt{\pi\mathcal{L}}\cdot C_{0}C_{1}^{2}\cdot R\cdot\sqrt{\frac{r\,d_{\text{out}}}{m}}.$
@@
For large $m$, the slower-decaying $O(1/\sqrt{m})$ concentration term dominates both bounds, and the comparison reduces to constants. For moderate $m$, the different structural dependences on architecture and data can make either bound tighter.
[…]
@@ E.9.1 Complementary Strengths
    Posterior concentration: When the posterior is highly concentrated around a good solution, $C_{\max}$ can be very small, making the KL term tight. In the extreme case of a near-deterministic posterior, the PAC-Bayes gap approaches zero while the Gaussian complexity bound remains fixed by the class geometry.
@@
    Large input norms: PAC-Bayes bounds do not depend on $\|X\|_{F}$, which is advantageous when inputs have very large norm or heavy tails. The Gaussian complexity bound’s leading term grows linearly in $\|X\|_{F}$.
@@
    High-dimensional inputs with moderate rank: Looking at the full expressions, PAC-Bayes scales with $r(d_{\text{in}}+d_{\text{out}})$ while Gaussian complexity (with bounded inputs) scales with $C_{0}^{2}C_{1}^{4}R^{2}\cdot r\,d_{\text{out}}$. When $d_{\text{in}}\gg d_{\text{out}}$ and $C_{0}^{2}C_{1}^{4}R^{2}$ is large, PAC-Bayes can still be favorable despite the $d_{\text{in}}$ term, because it avoids the spectral norm constants entirely. Concretely, PAC-Bayes is structurally favorable when $C_{\max}(d_{\text{in}}+d_{\text{out}})\lesssim\pi\mathcal{L}\,C_{0}^{2}C_{1}^{4}R^{2}\,d_{\text{out}}$, i.e., when the posterior is sufficiently concentrated relative to the spectral norm and Lipschitz constants.
Remark E.24 (When Gaussian complexity is stronger).
Gaussian complexity bounds can be tighter when:
  1. 1.
    Diffuse posteriors with bounded spectral norms: When $C_{\max}$ is large (the posterior has not concentrated tightly), the PAC-Bayes KL term can be vacuous. The Gaussian complexity bound depends on the spectral norm constraints $C_{1}$, not on posterior concentration, and can remain informative even when the posterior is spread over a large region of parameter space.
  2. 2.
    Normalized inputs with small spectral norm constants: When $\|x_{j}\|_{2}\leq R$ with $R$ moderate and spectral norms are well-controlled ($C_{1}$ not too large), the Gaussian complexity bound’s leading constant $\sqrt{\pi\mathcal{L}}\,C_{0}C_{1}^{2}R$ can be smaller than $\sqrt{C_{\max}/2}$. This is the regime where the Gaussian complexity bound is numerically tighter.
  3. 3.
    No prior specification required: Unlike PAC-Bayes, which requires choosing a prior $P$ and where the bound quality depends sensitively on this choice, Gaussian complexity bounds depend only on the realized network architecture and data. This makes them valuable as a prior-free sanity check.
@@
    Architectural transparency: The bound explicitly shows how rank, spectral norms, and layer dimensions interact through the network depth, providing guidance for architectural design that is not available from the PAC-Bayes bound’s KL-divergence summary.
@@ E.10 Alternative Approaches for High-Probability Spectral Norm Bounds
Both approaches require an additional regularity condition on the loss function:
Assumption E.27 (Convex loss).
The loss function $\ell(\cdot,y):\mathbb{R}^{k}\to[0,1]$ is convex in its first argument for every $y$.
Remark E.28.
Assumption E.27 is satisfied by most standard loss functions used in practice, including the squared loss $\ell(\hat{y},y)=\|\hat{y}-y\|^{2}$ (after rescaling to $[0,1]$), the cross-entropy loss (when composed with the softmax, the loss is convex in the logit vector), and the hinge loss. The convexity requirement is needed because the BNN predictor $f_{\mathrm{BNN}}$ is a mixture of $f_{\mathrm{BNN}}^{\mathrm{good}}$ and $f_{\mathrm{BNN}}^{\mathrm{bad}}$, and we need Jensen’s inequality to bound the loss of the mixture in terms of the losses of its components.
@@ E.10.1 Setup and Notation
Define the “good event” $\mathcal{E}_{\text{good}}$ as the intersection of all these events across all $D$ layers:
@@
We write $p:=\mathbb{P}(\mathcal{E}_{\text{good}})\geq 1-\delta_{\text{total}}$ throughout this section.
@@ E.10.2 Approach A: Decomposition of the BNN Predictor
By the law of total expectation applied to the posterior distribution $q$, we can decompose this as:
$f_{\mathrm{BNN}}(x;\theta)=p\cdot f_{\mathrm{BNN}}^{\text{good}}(x;\theta)+(1-p)\cdot f_{\mathrm{BNN}}^{\text{bad}}(x;\theta),$ (22)
where:
$f_{\mathrm{BNN}}^{\text{good}}(x;\theta)$ $:=\mathbb{E}_{(A,B)\sim q}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}],$ (23)
$f_{\mathrm{BNN}}^{\text{bad}}(x;\theta)$ $:=\mathbb{E}_{(A,B)\sim q}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}^{c}].$ (24)
Note: Both $f_{\mathrm{BNN}}^{\text{good}}$ and $f_{\mathrm{BNN}}^{\text{bad}}$ are deterministic functions of $x$ (given the variational parameters $\theta$). The conditioning in (23)–(24) is over the posterior $(A,B)\sim q$, not over the data distribution. Equation (22) is a pointwise identity for each $x$, expressing $f_{\mathrm{BNN}}(x)$ as a convex combination of two deterministic functions.
Analysis of the “good” term.
The conditional expectation $f_{\mathrm{BNN}}^{\text{good}}$ is taken over the conditional distribution $q(\cdot\mid\theta,\mathcal{E}_{\text{good}})$, which is supported on factor matrices satisfying the spectral norm bounds. By the same argument as Lemma E.11 (replacing $q$ by $q(\cdot\mid\mathcal{E}_{\text{good}})$):
@@
where $\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)$ consists of network functions realized by factor pairs in the support of $q(\cdot\mid\mathcal{E}_{\text{good}})$. Since all such pairs satisfy the spectral norm constraints by definition of $\mathcal{E}_{\text{good}}$:
@@
By the same arguments as Proposition E.13, the Gaussian complexity of the class containing $f_{\mathrm{BNN}}^{\text{good}}$ is bounded by that of Pinto’s class. Applying Theorem E.15, we obtain that with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$:
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}^{\text{good}}(x_{j}),y_{j})+\sqrt{\pi\mathcal{L}}\;\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}.$ (27)
From $f_{\mathrm{BNN}}^{\text{good}}$ to $f_{\mathrm{BNN}}$: applying loss convexity.
We now bound the population loss of the actual deployed predictor $f_{\mathrm{BNN}}$. By (22), for each $(x,y)$:
$f_{\mathrm{BNN}}(x)=p\cdot f_{\mathrm{BNN}}^{\text{good}}(x)+(1-p)\cdot f_{\mathrm{BNN}}^{\text{bad}}(x).$
Under Assumption E.27, $\ell(\cdot,y)$ is convex in its first argument, so by Jensen’s inequality:
$\ell(f_{\mathrm{BNN}}(x),y)$ $\leq p\cdot\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)+(1-p)\cdot\ell(f_{\mathrm{BNN}}^{\text{bad}}(x),y)$
$\leq p\cdot\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)+(1-p)\cdot 1,$ (28)
where in the second inequality we used $\ell\in[0,1]$.
Taking the expectation over $(x,y)\sim\mathcal{D}$:
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]\leq p\cdot\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+(1-p)\leq\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+\delta_{\text{total}},$ (29)
where the last step uses $1-p\leq\delta_{\text{total}}$ and $p\leq 1$.
@@
Substituting (27) into (29):
Proposition E.29 (Generalization bound under high-probability spectral norm constraints).
Let $\ell(\cdot,y)\in[0,1]$ be $\mathcal{L}$-Lipschitz and convex in its first argument (Assumptions in Theorem E.15 and Assumption E.27). Let high-probability bounds (17)–(18) hold with $\delta_{\text{total}}=\sum_{i=1}^{D}(\delta_{i}^{A}+\delta_{i}^{B})$. Then with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$:
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}^{\mathrm{good}}(x_{j}),y_{j})$
$\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}+\delta_{\text{total}},$ (30)
where $C_{i}=C_{i}^{A}C_{i}^{B}$ and $f_{\mathrm{BNN}}^{\mathrm{good}}$ is defined in (23).
Remark E.30 (Empirical loss of $f_{\mathrm{BNN}}^{\mathrm{good}}$ vs. $f_{\mathrm{BNN}}$).
The empirical loss term involves $f_{\mathrm{BNN}}^{\mathrm{good}}$ rather than $f_{\mathrm{BNN}}$, because the generalization bound from Theorem E.15 is applied to $f_{\mathrm{BNN}}^{\mathrm{good}}$ (which belongs to the controlled function class). In practice, $f_{\mathrm{BNN}}^{\mathrm{good}}$ can be computed via Monte Carlo sampling from the truncated posterior $q(\cdot\mid\mathcal{E}_{\text{good}})$, i.e., by rejection sampling: draw $(A,B)\sim q$ and retain only samples satisfying the spectral norm bounds. When $\delta_{\text{total}}$ is small, the rejection rate is low and $f_{\mathrm{BNN}}^{\mathrm{good}}\approx f_{\mathrm{BNN}}$.
@@
The additional term $\delta_{\text{total}}$ represents the “cost” of using high-probability bounds instead of almost-sure bounds. For Gaussian posteriors with Van Handel-type bounds (see Remark E.35), choosing $\delta_{i}^{A}=\delta_{i}^{B}=\delta_{0}/(2D)$ for each layer gives $\delta_{\text{total}}=\delta_{0}$. The spectral norm bounds $C_{i}^{A}$ and $C_{i}^{B}$ then scale as $O(\sqrt{\log(D/\delta_{0})})$, which introduces a mild logarithmic dependence on the depth and confidence level.
Remark E.32 (Why convexity is necessary).
Without Assumption E.27, the bound (28) does not hold. Using only Lipschitz continuity gives
$|\ell(f_{\mathrm{BNN}}(x),y)-\ell(f_{\mathrm{BNN}}^{\mathrm{good}}(x),y)|\leq\mathcal{L}\|f_{\mathrm{BNN}}(x)-f_{\mathrm{BNN}}^{\mathrm{good}}(x)\|_{2}=\mathcal{L}(1-p)\|f_{\mathrm{BNN}}^{\mathrm{bad}}(x)-f_{\mathrm{BNN}}^{\mathrm{good}}(x)\|_{2},$
which requires bounding $\|f_{\mathrm{BNN}}^{\mathrm{bad}}(x)\|_{2}$—the expected output under the “bad” event where spectral norm constraints are violated. This quantity can be arbitrarily large for unbounded variational families, making the Lipschitz approach impractical without additional assumptions on the tails of $q$. Loss convexity avoids this issue entirely by exploiting the $[0,1]$-boundedness of $\ell$.